evaluate f(x)=2x and g(x)=x+3
1. f(g(2))
2. g(f(2))

Answer by Jk22(389)


f(g(x))=2*g(x)=2*(x+3), f(g(2))=2*(2+3)=10
g(f(x))=2x+3, g(f(2))=2*2+3=7